Of the 24 seniors on the boy's soccer team 8 are taking History, 10 are taking Science, and 6 are taking both. How many are taking neither?
Found 2 solutions by richwmiller, checkley77:
Answer by richwmiller(9144)

24 total
2 history only
4 science only
6 both
12 one or both
12 neither

Answer by checkley77(12569)

24-(8+10)=24-18=6 STUDENTS ARE TAKING NEITHER.
NOTE: THE 6 TAKING BOTH ARE INCLUDED IN THE 18.